# Information theory for data-driven model reduction in physics and biology

### v3

@@ VII Chaotic systems
(a) Fractal dimension of the embedded Kuramoto-Sivashinsky attractor with varying encoding dimension. We estimate fractal dimension here by computing the correlation dimension [125, 139]. The gray dashed line shows the upper bound on the fractal dimension given by the Kaplan-Yorke conjecture [140, 141]. Green dashed line denotes the dimension of the attractor approximated for the full system, which lives in an $\mathbb{R}^{N}$, $N=128$ discretization of the system’s domain. b: Fractal dimension of the embedded system as a function of encoding dimension. Dashed line shows the fractal dimension of the system measured using the full state. c: Mutual information between the embedding and the future state, as a function of encoding dimension. d: At sufficiently high embedding dimension, encodings preserve the system’s Lyapunov exponent. Dashed line shows the true exponent of $\lambda\approx 0.048$. Inset shows the distance $\lvert\delta h\rvert$ as a function of time for points which start close together (initial distance is normalized to one for comparison).
VIII Cyanobacteria experiments
@@
Figure S15: Discovering slow collective variables in cyanobacteria populations (a) Fluorescent images of cyanobacteria colonies labeled with EYFP driven by the kaiBC promoter, allowing the visualization of Kai protein transcription. The colonies are imaged as they undergo cell growth and oscillations in Kai expression associated with the circadian rhythm. (b) Time series of one cropped section of an individual colony (see SI Fig. S9 for details). (c) The “state” used for variational IB is the time-lagged intensity field with lag time $\tau$. (d) Variational IB embeddings of time-lagged images into two dimensions (see Supplementary Movie 2). Every line corresponds to one colony’s evolution in time. Note the apparent oscillations of different radii. (e) Embedding into three dimensions. We orient our axes to correspond to the three principal components of the data. Thus, the projection onto the $h_{2}=0$ plane corresponds to a projection along the dominant two principal components. In this subspace, we see a similar structure as in the 2D embedding. (f) Mutual information between the future state and the encoding, given by $I(X_{t+\Delta t},H_{t})$, for varying dimension of the latent space. Increasing the dimension of the embedding space beyond two leads only to marginal increases in $I(X_{t+\Delta t},H_{t})$; this tells us “when to stop.” Each small point represents one training instance of the variational IB model, while the large point shows the maximum estimated value. Because the InfoNCE estimator is a lower bound on the true information, we consider the maximum as the estimated mutual information. The value of $I_{\max}=I(X_{t},X_{t+\Delta t})$ is the mutual information estimated for the true dynamics. (g) Selected trajectories in latent space with large and small radii. (h) Time-average latent radius of all colonies. (i) Mean synchronization order parameter of the intensity images; see SI for computation details. (j) Mean radius versus synchronization parameter for each colony. VIB identifies clusters of cells characterized by high (blue) or low (black) synchronization. As revealed to us after our analysis, this clustering corresponds to differing theophylline concentrations across experiments. Within each experimental movie (each of which contains 2-3 colonies) the radii are mostly constant. (k) Sample time series of weakly (black) and highly (blue) synchronized colonies. We apply a slight Gaussian blur to better visualize the bacteria boundaries.
IX Simulation Parameters
@@
The fluid flow simulations simulations are contained in the “Cylinder in Crossflow” dataset, downloaded from Ref. [142]. We take the dataset at Reynolds number 150, and interpolate the velocity field from the unstructured 6569-node mesh to a regular grid of $300\times 150$ pixels. The VIB networks are trained with $\beta=10^{7}$ and a time buffer $\Delta t$ chosen randomly between $[4,24]$ (see our comment on time randomization in “VIB for deterministic dynamics” in Section V).
@@
X Comparison of various dimensionality reduction methods
Variational IB (VIB) is by no means the only numerical method for performing data-driven model reduction. In this section we compare the performance of variational IB (VIB) to several common data-driven model reduction or inference methods. VIB uses a neural network to identify the relevant variables that are most predictive of the future. This construction learns dynamically relevant variables in contrast to methods such as principal component analysis or diffusion maps (in the absence of additional stringent assumptions), and it finds potentially non-linear variables in an agnostic way without requiring a tailored library of non-linearities (such as extended dynamic mode decomposition). In addition to these factors, VIB produces physically well-defined relevant variables – transfer operator eigenfunctions – in contrast to other deep methods such as (variational) autoencoders. In Fig. S17 we compare the low-dimensional latent space trajectories of the reduced models obtained using several of these standard methods to those output by VIB. In the remainder of this section we discuss these methods in detail.
@@ X Comparison of various dimensionality reduction methods
Figure S17: Comparison of various dimensionality reduction techniques on cyanobacteria dynamics Top row illustrates linear methods, bottom row shows non-linear methods. For each method, we plot the trajectories of all cyanobacteria colonies when projected onto the dominant modes (whose definition depends on the particular method). Blue trajectories show the highly synchronized colonies corresponding to high theophylline concentration, black/gray trajectories show the unsynchronized colonies from low theophylline conditions. Unlike the first four methods, deep neural networks (bottom right) directly produce a two-dimensional latent variable and do not provide access to subleading modes. (a) Principal component analysis reduces the dimensionality of data by finding a projection onto the directions of most variance (sketch). Here we show the projection of the cyanobacteria trajectories onto several pairs of principal components, $a_{i}(t)=\mathbf{x}(t)\cdot\mathbf{m}_{i}$. Each curve corresponds to the trajectory of one single colony evolving under one of four theophylline conditions; there are 10 trajectories in total, coming from 4 experimental conditions. Below, the first four principal components are shown. Here, the state is composed of two time-lagged images of the cyanobacteria colony; the top half of $\mathbf{m}$ corresponds to the earlier image, the bottom half corresponds to the lagged image. (b) Dynamic mode decomposition is performed on pairs of data points and aims to find a linear evolution operator $\mathbf{A}$ that links them (sketch). (Right) Trajectories of the projection of the state onto several eigenvectors of $\mathbf{A}$. The eigenvectors themselves are shown below (the imaginary parts of the first two modes are uniformly zero). (c) Independent component analysis seeks a linear projection of the data onto statistically independent components $\mathbf{s}_{i}$. In other words, the distribution $p(s_{i}|s_{j})$ is independent of $s_{j}$ (sketch). (Right, top) Trajectories of the projection of the state onto several independent components. (Bottom) columns of the mixing matrix; the components $s_{i}$ describe how the weighting of these columns evolves in time. (d) (Left) Diffusion maps build a graph of the observed state variables with edges weights determined by their distance from one another. (Right) Projection of the trajectories onto the first non-trivial eigenfunctions (because $a_{0}=$const) of the inferred transfer operator on the graph. (e) A (time-lagged) autoencoder is a deep neural network architecture which encodes the state $\mathbf{x}_{t}$ into a low-dimensional latent space. From the encoding one then tries to reconstruct the future state. (Right) Trajectories of the cyanobacterial colonies when encoded into a two-dimensional latent space. (f) A variational autoencoder is similar to a standard autoencoder, where one instead learns a distribution over possible encodings, $p(\mathbf{h}_{t}|\mathbf{x}_{t})$. (Right) Trajectories of the cyanobacterial colonies when encoded into a two-dimensional latent space. (g) Implementation of the network in Ref. [28] which encodes the state using a time-lagged autoencoder with an additional loss term that aims to enforce linearity of the embedded dynamics, $\mathbf{h}_{t+\Delta t}\approx\mathbf{K}\mathbf{h}_{t}$. (h) Latent trajectories produced by variational IB (VIB).
@@
The leveraging of non-Gaussianity in ICA is required due to the fact that, if using whitened data ($\mathbf{z}=\mathbf{V}\mathbf{x}$ so that $\mathbf{z}$ is zero mean $\mathbb{E}[\mathbf{z}]=\mathbf{0}$, and unit variance, $\mathbb{E}[\mathbf{z}\mathbf{z}^{T}]=\mathbf{I}$), then the mixing matrix $\mathbf{A}_{\text{ICA}}$ cannot be estimated if the components $\mathbf{s}$ are normal Gaussian variables: any orthogonal matrix will satisfy (S112). For dynamical systems, however, one can get around this requirement of non-Gaussianity. In particular, one can use a time-lagged variable $\mathbf{z}_{t-\tau}=\mathbf{A}_{\text{ICA}}\mathbf{s}_{t-\tau}$ to compute the mixing matrix from the time-correlation matrix of the signal $\mathbf{z}_{t}$ [24]. In particular, we have
@@
The VIB loss function is very similar to the $\beta$-VAE loss [149]. Rather than attempting to reconstruct the original state $x$, VIB replaces this term with an estimate of the mutual information between the latent variable $z$ and some other relevance variable $y$ (for us, $y=x_{t+\Delta t}$). In contrast to DMD and diffusion maps, neither VIB nor $\beta$-VAEs make any mention of transfer operators and are instead motivated by purely statistical considerations. As we show in the main text, the latent variables learned by VIB correspond to eigenfunctions of the transfer operator. While we cannot claim that that $\beta$-VAEs learn the same thing, some preliminary results in Fig. S18 suggest they may coincide to some degree. In the cyanobacteria dataset, we see that the latent variables learned by a $\beta$-VAE show the same qualitative structure as in VIB, but are less smooth (Fig. S17).
@@
As another example, [28] attempts to learn the full Koopman operator using neural networks. This can be thought of as an extension of eDMD, where instead of prescribing the library of nonlinear terms by hand, they can be learned by a neural network. The latent dynamics are then encouraged to be linear by minimizing the deviation between the true future (latent) state $\mathbf{z}_{t+\Delta t}$ and its linear approximation found by least squares. Similar approaches have also been explored in Refs. [30, 29]. As an illustration, we trained one such network on the cyanobacteria which can be seen in Fig. S17. This approach was applied primarily to deterministic systems. We expect that combining their method of encouraging linear latent dynamics together with the VIB objective function may be fruitful and lead to more well-behaved latent variables.
Appendix A Appendix: Details for the calculation of the Hessian of the IB Lagrangian
First order derivatives for $n\neq 0$ are given by
$\partial_{n}^{\mu}p(h_{\nu}|x)\equiv\frac{\partial}{\partial f_{n}(h_{\mu})}p(h_{\nu}|x)$ $=p(h_{\nu}|x)\delta_{\mu\nu}\tilde{\phi}_{n}(x)-\frac{f_{0}^{\nu}\exp\left[\sum_{n}\tilde{\phi}_{n}(x)f_{n}^{\nu}\right]}{\left(\sum_{h_{\lambda}}f_{0}^{\lambda}\exp\left[\sum_{n}\tilde{\phi}_{n}(x)f_{n}^{\lambda}\right]\right)^{2}}\left[\sum_{h_{\lambda}}f_{0}^{\lambda}\exp\left[\sum_{n}\tilde{\phi}_{n}(x)f_{n}^{\lambda}\right]\tilde{\phi}_{n}\delta_{\mu\lambda}\right]$ (S113)
$=p(h_{\nu}|x)\tilde{\phi}_{n}(x)\left(\delta_{\mu\nu}-p(h_{\mu}|x)\right)$ (S114)
and for $n=0$,
$\partial_{n}^{\mu}p(h_{\nu}|x)$ $=p(h_{\nu}|x)\frac{1}{f_{0}^{\mu}}\left(\delta_{\mu\nu}-p(h_{\mu}|x)\right).$ (S115)
We assign $\tilde{\phi}_{0}=\frac{1}{f_{0}^{\mu}}$, taking care to track the correct Greek index which otherwise is not present for $\tilde{\phi}_{n}$. Subsequently,
$\partial_{n}^{\mu}p(h_{\nu})$ $=\int\text{d}{x}\,p(x)p(h_{\nu}|x)\tilde{\phi}_{n}(x)\left(\delta_{\mu\nu}-p(h_{\mu}|x)\right)$ (S116)
As it turns out, we do not need second-order derivatives of $p(h_{\nu}|x)$, as only first-order terms appear in the Hessian. In the following, for brevity we denote the marginals simply by $p(h_{\lambda})=p_{\lambda}$.
We are now in a position to calculate the Hessian of the Lagrangian. We begin with the compression term,
$\partial_{n}^{\mu}\partial_{m}^{\nu}I(X,H)$ $=\partial_{n}^{\mu}\partial_{m}^{\nu}\int\text{d}{x}\sum_{h_{\lambda}}p(h_{\lambda}|x)p(x)\log\left(\frac{p(h_{\lambda}|x)}{p(h_{\lambda})}\right)$
$=\int\text{d}x\sum_{h_{\lambda}}p(x)(\partial_{n}^{\mu}\partial_{m}^{\nu}p(h_{\lambda}|x))\log\left(\frac{p(h_{\lambda}|x)}{p(h_{\lambda})}\right)+p(x)\partial_{m}^{\nu}p(h_{\lambda}|x)\frac{\partial_{n}^{\mu}p(h_{\lambda}|x)}{p(h_{\lambda}|x)}-p(x)\partial_{m}^{\nu}p(h_{\lambda}|x)\frac{\partial_{n}^{\mu}p(h_{\lambda})}{p(h_{\lambda})}$
$\quad+\int\text{d}x\sum_{h_{\lambda}}p(x)\partial_{n}^{\mu}p(h_{\lambda}|x)\frac{\partial_{m}^{\nu}p(h_{\lambda}|x)}{p(h_{\lambda}|x)}+p(x)p(h_{\lambda}|x)\frac{\partial_{n}^{\mu}\partial_{m}^{\nu}p(h_{\lambda}|x)}{p(h_{\lambda}|x)}-p(x)p(h_{\lambda}|x)\frac{\partial_{n}^{\mu}p(h_{\lambda}|x)}{p(h_{\lambda}|x)^{2}}\partial_{m}^{\nu}p(h_{\lambda}|x)$
$\quad+\int\text{d}x\sum_{h_{\lambda}}p(x)\partial_{n}^{\mu}p(h_{\lambda}|x)\frac{\partial_{m}^{\nu}p_{\lambda}}{p_{\lambda}}+p(x)p(h_{\lambda}|x)\frac{\partial_{n}^{\mu}\partial_{m}^{\nu}p_{\lambda}}{p_{\lambda}}-p(x)p(h_{\lambda}|x)\frac{\partial_{n}^{\mu}p_{\lambda}}{p_{\lambda}^{2}}\partial_{m}^{\nu}p_{\lambda}$
by evaluating at the uniform encoder, many terms drop out
$=\int\text{d}x\sum_{h_{\lambda}}\frac{p(x)}{p_{\lambda}}\left[\partial_{m}^{\nu}p(h_{\lambda}|x)\partial_{n}^{\mu}p(h_{\lambda}|x)-\partial_{m}^{\nu}p(h_{\lambda}|x)\partial_{n}^{\mu}p(h_{\lambda})\right]$
$\quad+\int\text{d}x\sum_{h_{\lambda}}\frac{p(x)}{p_{\lambda}}\left[\partial_{n}^{\mu}p(h_{\lambda}|x)\partial_{m}^{\nu}p_{\lambda}-\partial_{n}^{\mu}p_{\lambda}\partial_{m}^{\nu}p_{\lambda}\right]$
$=\int\text{d}x\sum_{h_{\lambda}}\frac{p(x)}{p_{\lambda}}\partial_{m}^{\nu}p(h_{\lambda}|x)\partial_{n}^{\mu}p(h_{\lambda}|x)-\sum_{h_{\lambda}}\frac{1}{p_{\lambda}}\partial_{m}^{\nu}p_{\lambda}\partial_{n}^{\mu}p_{\lambda}$
where for the vanishing second line we used that $\partial_{n}^{\mu}\sum_{h_{\lambda}}p(h_{\lambda}|x)=0$ (and analogously for terms with a single $p_{\lambda}$), as well as the fact that $\sum_{x}p(x)\partial_{n}^{\mu}p(h_{\lambda}|x)=\partial_{n}^{\nu}p_{\lambda}$. We now substitute the expressions for the derivatives, evaluated at the uniform encoder,
$\partial_{n}^{\mu}\partial_{m}^{\nu}I(X,H)$ $=\sum_{x,h_{\lambda}}\frac{p(x)}{p_{\lambda}}(p_{\lambda}\tilde{\phi}_{m}(x)\left(\delta_{\lambda\nu}-p_{\nu}\right))(p_{\lambda}\tilde{\phi}_{n}(x)\left(\delta_{\lambda\mu}-p_{\mu}\right))$
$\quad-\sum_{h_{\lambda}}\frac{1}{p_{\lambda}}(p_{\lambda}\langle\tilde{\phi}_{m}(x)\rangle\left(\delta_{\lambda\nu}-p_{\nu}\right))(p_{\lambda}\langle\tilde{\phi}_{n}(x)\rangle\left(\delta_{\lambda\mu}-p_{\mu}\right))$
$=\left(\langle\tilde{\phi}_{n}\tilde{\phi}_{m}\rangle-\langle\tilde{\phi}_{n}\rangle\langle\tilde{\phi}_{m}\rangle\right)\sum_{h_{\lambda}}p_{\lambda}\left(\delta_{\lambda\nu}\delta_{\lambda\mu}-p_{\nu}\delta_{\lambda\mu}-p_{\mu}\delta_{\lambda\nu}+p_{\nu}p_{\mu}\right)$
$=\left(\langle\tilde{\phi}_{n}\tilde{\phi}_{m}\rangle-\langle\tilde{\phi}_{n}\rangle\langle\tilde{\phi}_{m}\rangle\right)\left(\delta_{\mu\nu}p_{\mu}-p_{\nu}p_{\mu}\right)$ (S117)
We now perform the same calculation for the prediction term,
$\partial_{n}^{\mu}\partial_{m}^{\nu}I(Y,H)$ $=\partial_{n}^{\mu}\partial_{m}^{\nu}\sum_{y,h_{\lambda}}p(y,h_{\lambda})\log\left(\frac{p(y,h_{\lambda})}{p_{\lambda}p(y)}\right)$
$=\int\text{d}y\sum_{h_{\lambda}}\partial_{m}^{\nu}p(y,h_{\lambda})\left(\frac{\partial_{n}^{\mu}p(y,h_{\lambda})}{p(y,h_{\lambda})}-\frac{\partial_{n}^{\mu}p_{\lambda}}{p_{\lambda}}\right)+\partial_{n}^{\mu}\partial_{m}^{\nu}p(y,h_{\lambda})$
$=\int\text{d}y\sum_{h_{\lambda}}\frac{1}{p(y)p_{\lambda}}\partial_{m}^{\nu}p(y,h_{\lambda})\partial_{n}^{\mu}p(y,h_{\lambda})-\sum_{h_{\lambda}}\frac{1}{p_{\lambda}}\partial_{m}^{\nu}p_{\lambda}\partial_{n}^{\mu}p_{\lambda}$
$=\int\text{d}y\sum_{h_{\lambda}}\frac{1}{p(y)p_{\lambda}}\left(\int\text{d}xp(y,x)\partial_{m}^{\nu}p(h_{\lambda}|x)\right)\left(\int\text{d}x^{\prime}p(y,x^{\prime})\partial_{n}^{\mu}p(h_{\lambda}|x^{\prime})\right)-\sum_{h_{\lambda}}\frac{1}{p_{\lambda}}\partial_{m}^{\nu}p_{\lambda}\partial_{n}^{\mu}p_{\lambda}$
$=\sum_{y,h_{\lambda}}\frac{1}{p(y)p_{\lambda}}\left(\int\text{d}xp(y,x)p_{\lambda}\tilde{\phi}_{m}(x)(\delta_{\lambda\nu}-p_{\nu})\right)\left(\int\text{d}x^{\prime}p(y,x^{\prime})p_{\lambda}\tilde{\phi}_{n}(x^{\prime})(\delta_{\lambda\mu}-p_{\mu})\right)-\sum_{h_{\lambda}}\frac{1}{p_{\lambda}}\partial_{m}^{\nu}p_{\lambda}\partial_{n}^{\mu}p_{\lambda}$
$=\left(\left\langle\frac{\langle\tilde{\phi}_{n}\rangle_{p(\cdot,y)}\langle\tilde{\phi}_{m}\rangle_{p(\cdot,y)}}{p(y)}\right\rangle_{y}-\langle\tilde{\phi}_{n}\rangle\langle\tilde{\phi}_{m}\rangle\right)\left(\delta_{\mu\nu}p_{\mu}-p_{\nu}p_{\mu}\right)$
Appendix B Appendix: Optimal encoding for Hopf normal form dynamics
$\dot{x}$ $=(\mu-x^{2}-y^{2})x-\omega y$ (S118)
$\dot{y}$ $=(\mu-x^{2}-y^{2})y-\omega x$ (S119)
Expressed in polar coordinates, we have
$=r(\mu-r^{2})(\partial_{r}\phi)+\omega\partial_{\theta}\phi$ (S123)
For the angle coordinate alone the situation is much simpler,
$\omega\partial_{\theta}\phi(\theta)=\lambda\phi(\theta)$
$\phi_{n}(\theta)=e^{in\theta},\quad\quad\lambda_{n}=in\omega.$
We next ask whether an encoder of the form $p(h|\theta)\propto\delta(h-\theta)$ is a solution. Recall that the probability distribution appearing in (S126) is a distribution over future positions, $p(R_{t+\Delta t}=r^{\prime},\Theta_{t+\Delta t}=\theta+\omega\Delta t|h)$. This distribution can be calculated as
$\propto\delta((\theta_{t+\Delta t}-\omega\Delta t)-h)$
where we used that the two terms in the integrand of the first equation are both delta functions. Plugging this into (S126) shows that $p(h|\theta)\propto\delta(h-\theta)$ is consistent, and hence is an optimal encoding. In the presence of noise, the eigenvectors are perturbed and may gain a dependence on $r$ (see Fig. S19).
Appendix C Appendix: Stability analysis of $I(H_{t},H_{t+\Delta t})$
$=p(h_{\nu})(\delta_{\mu\nu}-p(h_{\mu}))\delta_{n0}$ (S128)
$=\int\text{d}x_{0}\text{d}x_{t}\,p(h_{t}|x_{t})p(h_{0}|x_{0})p(x_{0},x_{t})$
Evaluated at the uniform encoder, the first term disappears because we have the derivative of a constant. From the equations for the derivatives (S127) and (S128), we can conclude $\partial_{n}^{\nu}\partial_{m}^{\mu}I(H_{0},H_{t})\propto\delta_{n0}\delta_{m0}$, which shows that we lose all information about the eigenfunctions. The Hessian of the Lagrangian for IB modified by the term $I(H_{0},H_{t})$ will therefore have the same stability properties as the original IB problem.

### v4

@@ E.4 Summary
(a) Fractal dimension of the embedded Kuramoto-Sivashinsky attractor with varying encoding dimension. We estimate fractal dimension here by computing the correlation dimension [162, 176]. Green dashed line denotes the dimension of the attractor approximated for the full system, which lives in an $\mathbb{R}^{N}$, $N=128$ discretization of the system’s domain. b: Fractal dimension of the embedded system as a function of encoding dimension. Dashed line shows the fractal dimension of the system measured using the full state. c: Mutual information between the embedding and the future state, as a function of encoding dimension. d: At sufficiently high embedding dimension, encodings preserve the system’s Lyapunov exponent. Dashed line shows the true exponent of $\lambda\approx 0.048$. Inset shows the distance $\lvert\delta h\rvert$ as a function of time for points which start close together (initial distance is normalized to one for comparison).
S8 Cyanobacteria experiments
@@
S9 Simulation Parameters
@@
The fluid flow simulations simulations are contained in the “Cylinder in Crossflow” dataset, downloaded from Ref. [177]. We take the dataset at Reynolds number 150, and interpolate the velocity field from the unstructured 6569-node mesh to a regular grid of $300\times 150$ pixels. The VIB networks are trained with $\beta=10^{7}$ and a time buffer $\Delta t$ chosen randomly between $[4,24]$ (see our comment on time randomization in “VIB for deterministic dynamics” in Section S5).
@@
S10 Comparison of various dimensionality reduction methods
Variational IB (VIB) is by no means the only numerical method for performing data-driven model reduction. In this section we compare the performance of variational IB (VIB) to several common data-driven model reduction or inference methods. VIB uses a neural network to identify the relevant variables that are most predictive of the future. This construction learns dynamically relevant variables in contrast to methods such as principal component analysis or diffusion maps (in the absence of additional stringent assumptions), and it finds potentially non-linear variables in an agnostic way without requiring a tailored library of non-linearities (such as extended dynamic mode decomposition). In addition to these factors, VIB produces physically well-defined relevant variables – transfer operator eigenfunctions – in contrast to other deep methods such as (variational) autoencoders. In Fig. 8 we compare the low-dimensional latent space trajectories of the reduced models obtained using several of these standard methods to those output by VIB. In the remainder of this section we discuss these methods in detail.
@@
The leveraging of non-Gaussianity in ICA is required due to the fact that, if using whitened data ($\mathbf{z}=\mathbf{V}\mathbf{x}$ so that $\mathbf{z}$ is zero mean $\mathbb{E}[\mathbf{z}]=\mathbf{0}$, and unit variance, $\mathbb{E}[\mathbf{z}\mathbf{z}^{T}]=\mathbf{I}$), then the mixing matrix $\mathbf{A}_{\text{ICA}}$ cannot be estimated if the components $\mathbf{s}$ are normal Gaussian variables: any orthogonal matrix will satisfy Eq. (S132). For dynamical systems, however, one can get around this requirement of non-Gaussianity. In particular, one can use a time-lagged variable $\mathbf{z}_{t-\tau}=\mathbf{A}_{\text{ICA}}\mathbf{s}_{t-\tau}$ to compute the mixing matrix from the time-correlation matrix of the signal $\mathbf{z}_{t}$ [129]. In particular, we have
@@
The VIB loss function is very similar to the $\beta$-VAE loss [183]. Rather than attempting to reconstruct the original state $x$, VIB replaces this term with an estimate of the mutual information between the latent variable $z$ and some other relevance variable $y$ (for us, $y=x_{t+\Delta t}$). In contrast to DMD and diffusion maps, neither VIB nor $\beta$-VAEs make any mention of transfer operators and are instead motivated by purely statistical considerations. As we show in the main text, the latent variables learned by VIB correspond to eigenfunctions of the transfer operator. While we cannot claim that that $\beta$-VAEs learn the same thing, some preliminary results in Fig. S17 suggest they may coincide to some degree. In the cyanobacteria dataset, we see that the latent variables learned by a $\beta$-VAE show the same qualitative structure as in VIB, but are less smooth (Fig. 8).
@@
As another example, [49] attempts to learn the full Koopman operator using neural networks. This can be thought of as an extension of eDMD, where instead of prescribing the library of nonlinear terms by hand, they can be learned by a neural network. The latent dynamics are then encouraged to be linear by minimizing the deviation between the true future (latent) state $\mathbf{z}_{t+\Delta t}$ and its linear approximation found by least squares. Similar approaches have also been explored in Refs. [114, 48]. As an illustration, we trained one such network on the cyanobacteria which can be seen in Fig. 8. This approach was applied primarily to deterministic systems. We expect that combining their method of encouraging linear latent dynamics together with the VIB objective function may be fruitful and lead to more well-behaved latent variables.